# RINO: Accurate, Robust Radar-Inertial Odometry with Non-Iterative Estimation*

### v3

@@ Abstract
Odometry in adverse weather conditions, such as fog, rain, and snow, presents significant challenges, as traditional vision- and LiDAR-based methods often suffer from degraded performance. Radar-Inertial Odometry (RIO) has emerged as a promising solution due to its resilience in such environments. In this paper, we present RINO, a non-iterative RIO framework implemented in an adaptively loosely coupled manner. Building upon ORORA as the baseline for radar odometry, RINO introduces several key advancements, including improvements in keypoint extraction, motion distortion compensation, and pose estimation via an adaptive voting mechanism. This voting strategy facilitates efficient polynomial-time optimization while simultaneously quantifying the uncertainty in the radar module’s pose estimation. The estimated uncertainty is subsequently integrated into the maximum a posteriori (MAP) estimation within a Kalman filter framework. Unlike prior loosely coupled odometry systems, RINO not only retains the global and robust registration capabilities of the radar component but also dynamically accounts for the real-time operational state of each sensor during fusion. Experimental results conducted on publicly available datasets demonstrate that RINO reduces translation and rotation errors by 1.06% and 0.09°/100m, respectively, when compared to the baseline method, thus significantly enhancing its accuracy. Furthermore, RINO achieves performance comparable to state-of-the-art methods.
@@ III-B State Estimations
Based on Eq.(2) and Eq.(3), we perform the prediction process of the ESKF. Since the noise has been incorporated into the error states, the prediction of nominal state variables can be directly obtained from equation Eq.(2). As for the error states,
@@ III-D Pointcloud Registration With Non-iterative Solving
$\boldsymbol{q}_{ij}$ $=\boldsymbol{R}^{\circ}\boldsymbol{p}_{ij}+\boldsymbol{o}_{ij}+\boldsymbol{\varepsilon}_{ij},i,j=1,...,N,i\neq j$
@@
$\|\boldsymbol{q}_{ij}\|$ $=\|\boldsymbol{p}_{ij}\|+\widetilde{o}_{ij}+\widetilde{\varepsilon}_{ij},i,j=1,...,N,i\neq j$
@@
$\theta_{k}=\mathrm{Ang}(\boldsymbol{q}_{k})-\mathrm{Ang}(\boldsymbol{p}_{k}),\ k=1,...,K,$ (20)
$\boldsymbol{t}_{i}=\boldsymbol{q}_{i}-\hat{\boldsymbol{R}}\boldsymbol{p}_{i},\ i=1,...,N,$ (21)
@@
$\hat{x}$ $=\mathop{\arg\min}\limits_{x\in\mathbb{R}}\sum_{m=1}^{M}\min\bigg{\{}\frac{1}{\sigma_{x,m}^{2}}(x-x_{m})^{2},\bar{c}^{2}\bigg{\}},m=1,...,M$ (24)
@@ IV-A1 Dataset
The real-world vehicle tests were carried out in a factory area located in the Shunyi District of Beijing. An experimental vehicle, equipped with a scanning radar, an inertial measurement unit (IMU), and real-time kinematic (RTK) positioning, was employed to gather data for these tests. We have designated these data as “Shunyi”. The primary aim of the real-world vehicle experiments was to validate the algorithm’s viability for deployment in authentic applications per step.
@@ IV-A4 Parameters
The number of parameters in RINO is small which reflects the model’s strong generalization and adaptability to new data. We listed all the parameters in the Tab.I. The z-value zw_ is used to threshold the noise, and wf_ is the median filter width; both are tunable parameters in the improved cen2018 keypoint extraction module. The meanings of min_for_mc_, max_for_mc_, and thr_for_stop_ have been explained earlier in the text. thr_for_inlier_ is the threshold used in MCIS to determine whether a point is an inlier. The parameters cbar_radial_ and cbar_tangential_ correspond to the values of $\bar{c}^{2}$ in Eq.(16) and Eq.(17). They are used for testing on the MulRan dataset and those set for real-world tests. The differences between the two sets of parameters are primarily due to the use of a different scanning radar type. It is important to note that the parameters suitable for the MulRan were obtained through adjustments on the DCC01 sequence and were kept fixed during testing on the remaining sequences, as We believe that this approach to renders the experimental results more meaningful.
@@
SuMa 2.71/0.4 4.07/0.9 2.14/0.6 2.90/0.8 2.64/0.6 2.17/0.6 1.66/0.6 1.49/0.5 1.65/0.4
LOAM 3.16/0.86 2.64/0.74 2.23/0.74 2.70/0.82 2.80/0.84 7.54/0.79 4.25/0.88 4.14/0.90 4.21/1.02
CFEAR-1 2.73/0.73 1.82/0.60 1.77/0.62 2.62/0.97 2.45/0.90 2.85/1.08 2.55/0.90 2.71/0.82 3.56/0.82
CFEAR-2 2.44/0.63 1.65/0.54 1.41/0.50 2.12/0.81 1.93/0.74 2.08/0.87 2.30/0.80 2.07/0.66 2.60/0.59
CFEAR-3 2.28/0.54 1.49/0.46 1.47/0.48 1.59/0.66 1.62/0.66 1.73/0.78 1.59/0.83 1.39/0.51 1.41/0.40
CFEAR-3-s50 2.09/0.55 1.38/0.47 1.26/0.47 1.48/0.65 1.51/0.63 1.59/0.75 1.62/0.62 1.35/0.52 1.19/0.37
SDPO 1.55/0.35 1.53/0.33 1.60/0.30 1.57/0.29 1.61/0.35 1.59/0.32 1.61/0.26 1.59/0.27 1.62/0.29
$\text{R}^{3}\text{O}$ 2.39/0.43 1.40/0.34 1.48/0.41 1.89/0.63 1.55/0.53 1.53/0.50 1.34/0.39 1.98/0.53 1.81/0.57
MC-RANSAC 4.79/1.27 3.76/1.00 4.39/1.40 - 6.60/1.74 4.67/1.19 4.76/1.08 6.49/1.73 7.72/2.19
MC-RANSAC + DPLR 4.49/1.12 3.64/1.00 4.37/1.39 - 6.32/1.68 4.42/1.12 5.92/1.43 8.38/2.08 6.61/1.88
ORORA(Cen2018) 3.18/0.66 2.38/0.57 2.77/0.79 - 3.12/0.78 2.53/0.57 3.51/0.76 3.31/0.76 2.79/0.64
ORORA(Cen2019) 3.12/0.67 2.60/0.51 2.37/0.57 - 3.28/0.82 3.04/0.70 3.53/0.84 2.67/0.64 2.11/0.49
RINO(Ours) 1.64/0.46 1.25/0.30 0.91/0.39 1.15/0.43 1.42/0.46 2.08/1.08 2.28/0.53 2.35/0.80 2.28/0.47
@@ IV-B Benchmark Results Within MulRan Dataset
We selected the best odometry or SLAM algorithms using different sensor modalities in large-scale environments as benchmarks. These methods include LOAM [7], SuMa [41], and the state-of-the-art scanning radar odometry algorithms introduced in Sec.II: CFEAR [14], SDPO [15], $\text{R}^{3}\text{O}$ [16], MC-RANSAC [31], and ORORA [17].
@@
We conducted tests on nine sequences from the MulRan dataset: DCC01, DCC02, DCC03, KAIST01, KAIST02, KAIST03, Riverside01, Riverside02, and Riverside03. The detailed results are shown in Table II. In the table, the red data points represent the minimum errors, while the blue data points correspond to the second smallest errors. It is recommended to view the table in color for improved clarity. It is important to note that, due to the unavailability of direct reproductions of these methods, the data presented in the table is taken from the reported results in publicly published papers. The table categorizes the aforementioned methods into LiDAR odometry, state-of-the-art radar odometry methods, and the baselines. The data is presented in the format (%/°/100m), where the first value represents the relative translation error, and the second value represents the relative rotation error.
@@
The proposed method, RINO, demonstrates relatively small relative translation and rotation errors across the majority of sequences, with particularly impressive performance in the DCC and KAIST sequences. RINO significantly improves upon the baseline, ORORA. Specifically, compared to ORORA with Cen2018 and Cen2019 keypoint extraction, RINO achieves average reductions of 1.43% and 1.73% in translation error, and 0.20°/100m and 0.13°/100m in rotation error in the DCC and KAIST sequences, respectively. In the DCC03 sequence, the error reduction reaches up to 1.46% in translation and 0.28°/100m in rotation. Moreover, when compared to LiDAR-based methods such as LOAM and SuMa, which offer higher resolution and reduced motion distortion per frame, RINO has demonstrated superior performance. We conclude that RINO achieves radar odometry performance comparable to state-of-the-art methods.
@@
When RINO is tested on the Riverside sequences, there is a notable decline in odometry accuracy compared to the DCC and KAIST sequences. The Riverside sequences, which were collected along a riverbank, represent a degenerate scenario that challenges odometry systems. Furthermore, the performance degradation of RINO is more pronounced in these sequences compared to other algorithms. A possible explanation for this lies in the Markovian assumption inherent in our pipeline. Both the Extended Kalman Filter (ESKF) and the odometry based on global point cloud registration operate in an incremental fashion, where the current pose transformation estimate is solely dependent on the previous time step. This approach, to some extent, overlooks the broader structural information of point clouds in the local map, potentially leading to confusion in degenerate environments.
@@ IV-C1 Effectiveness of Adaptive Voting
We implemented a scanning radar odometry method using progressive non-convex optimization to solve pointcloud registration problem (RO(GNC)), which represents ORORA receiving sensor data via ROS (Robot Operating System). For comparison, we also implemented a scanning radar odometry method using adaptive voting to solve pointcloud registration (RO(AV)), which represents the scanning radar branch in RINO. Both RO(GNC) and RO(AV) employ the improved keypoint extraction technique described in Sec. III-C, which differs from the code released by ORORA. Before feature matching, no voxel grid filtering is applied to the pointcloud in either method. This decision was based on our observations during the replication process, where we found that performing this step disrupted the timestamps assigned sequentially to each keypoint, thereby invalidating the motion distortion compensation. RO(GNC) performance is degraded compared to the test results presented in the ORORA paper (Tab.II). This degradation is hypothesized to arise from differences in parameter settings and the loss of precise azimuth encoding information in the raw scanning radar data. The comparison results between the two methods are shown in Tab.III. The odometry errors of RO(GNC) and RO(AV) are similar across multiple sequences. However, in subsequent runtime tests (Fig.10), we found that the running time of RO(AV) is generally shorter than that of RO(GNC). Therefore, we conclude that the adaptive voting method improves the system’s efficiency.
@@ IV-C4 Visualization
IV-D Runtime
We measured the time taken by each component, from receiving the scanning radar data to completing the pose estimation within a single run. Similar to ORORA, the majority of the system’s runtime is spent on keypoint extraction and matching. The time consumed by point cloud registration primarily depends on the number of keypoint pairs involved in the coarse matching process. The system’s efficiency decreases significantly when the vehicle is moving slowly or stationary.
The average runtime of the system is approximately 0.1381s, which is faster than the scanning millimeter-wave radar data acquisition frequency (4Hz). By comparing Fig.10(a) with Fig.10(b), we observe that the efficiency of the point cloud registration module improves by nearly 11.5% due to the adoption of adaptive voting.
IV-E Discussion about Estimation Uncertainty

### v4

@@ Abstract
Odometry in adverse weather conditions, such as fog, rain, and snow, presents significant challenges, as traditional vision- and LiDAR-based methods often suffer from degraded performance. Radar-Inertial Odometry (RIO) has emerged as a promising solution due to its resilience in such environments. In this paper, we present RINO, a non-iterative RIO framework implemented in an adaptively loosely coupled manner. Building upon ORORA as the baseline for radar odometry, RINO introduces several key advancements, including improvements in keypoint extraction, motion distortion compensation, and pose estimation via an adaptive voting mechanism. This voting strategy facilitates efficient polynomial-time optimization while simultaneously quantifying the uncertainty in the radar module’s pose estimation. The estimated uncertainty is subsequently integrated into the maximum a posteriori (MAP) estimation within a Kalman filter framework. Unlike prior loosely coupled odometry systems, RINO not only retains the global and robust registration capabilities of the radar component but also dynamically accounts for the real-time operational state of each sensor during fusion. Experimental results conducted on publicly available datasets demonstrate that RINO reduces translation and rotation errors by 1.06% and 0.09°/100m, respectively, when compared to the baseline method, thus significantly enhancing its accuracy. Furthermore, RINO achieves performance comparable to state-of-the-art methods. Our code is available at https://github.com/yangsc4063/rino.
@@ III-B State Estimations
Based on Eq.(2) and Eq.(3), we perform the prediction process of the Error-State Kalman Filter (ESKF). Since the noise has been incorporated into the error states, the prediction of nominal state variables can be directly obtained from equation Eq.(2). As for the error states,
@@ III-D Pointcloud Registration With Non-iterative Solving
$\boldsymbol{q}_{ij}$ $=\boldsymbol{R}^{\circ}\boldsymbol{p}_{ij}+\boldsymbol{o}_{ij}+\boldsymbol{\varepsilon}_{ij},i,j=1,.,N,i\neq j$
@@
$\|\boldsymbol{q}_{ij}\|$ $=\|\boldsymbol{p}_{ij}\|+\widetilde{o}_{ij}+\widetilde{\varepsilon}_{ij},i,j=1,.,N,i\neq j$
@@
$\theta_{k}=\mathrm{Ang}(\boldsymbol{q}_{k})-\mathrm{Ang}(\boldsymbol{p}_{k}),\ k=1,.,K,$ (20)
$\boldsymbol{t}_{i}=\boldsymbol{q}_{i}-\hat{\boldsymbol{R}}\boldsymbol{p}_{i},\ i=1,.,N,$ (21)
@@
$\hat{x}$ $=\mathop{\arg\min}\limits_{x\in\mathbb{R}}\sum_{m=1}^{M}\min\bigg{\{}\frac{1}{\sigma_{x,m}^{2}}(x-x_{m})^{2},\bar{c}^{2}\bigg{\}},m=1,.,M$ (24)
@@ IV-A1 Dataset
The real-world vehicle tests were carried out in a factory area located in the Shunyi District of Beijing. An experimental vehicle, equipped with a scanning radar, an inertial measurement unit (IMU), and real-time kinematic (RTK) positioning, was employed to gather data for these tests. We have designated these data as “Shunyi”. The aim of the real-world vehicle experiments was to validate the algorithm’s viability for deployment in authentic applications.
@@ IV-A4 Parameters
The number of parameters in RINO is small which reflects the model’s strong generalization and adaptability to new data. We listed all the parameters in the Tab.I. The z-value zw_ is used to threshold the noise, and wf_ is the median filter width; both are tunable parameters in the improved cen2018 keypoint extraction module. The meanings of min_for_mc_, max_for_mc_, and thr_for_stop_ have been explained earlier in the text. thr_for_inlier_ is the threshold used in MCIS to determine whether a point is an inlier. The parameters cbar_radial_ and cbar_tangential_ correspond to the values of $\bar{c}^{2}$ in Eq.(16) and Eq.(17). The differences between the two sets of parameters are primarily due to the use of a different scanning radar type. It is important to note that the parameters suitable for the MulRan were obtained through adjustments on the DCC01 sequence and were kept fixed during testing on the remaining sequences, as We believe that this approach to renders the experimental results more meaningful.
@@
SuMa [41] 2.71/0.4 4.07/0.9 2.14/0.6 2.90/0.8 2.64/0.6 2.17/0.6 1.66/0.6 1.49/0.5 1.65/0.4
LOAM [7] 3.16/0.86 2.64/0.74 2.23/0.74 2.70/0.82 2.80/0.84 7.54/0.79 4.25/0.88 4.14/0.90 4.21/1.02
CFEAR-1 [14] 2.73/0.73 1.82/0.60 1.77/0.62 2.62/0.97 2.45/0.90 2.85/1.08 2.55/0.90 2.71/0.82 3.56/0.82
CFEAR-2 [14] 2.44/0.63 1.65/0.54 1.41/0.50 2.12/0.81 1.93/0.74 2.08/0.87 2.30/0.80 2.07/0.66 2.60/0.59
CFEAR-3 [14] 2.28/0.54 1.49/0.46 1.47/0.48 1.59/0.66 1.62/0.66 1.73/0.78 1.59/0.83 1.39/0.51 1.41/0.40
CFEAR-3-s50 [14] 2.09/0.55 1.38/0.47 1.26/0.47 1.48/0.65 1.51/0.63 1.59/0.75 1.62/0.62 1.35/0.52 1.19/0.37
SDPO [15] 1.55/0.35 1.53/0.33 1.60/0.30 1.57/0.29 1.61/0.35 1.59/0.32 1.61/0.26 1.59/0.27 1.62/0.29
$\text{R}^{3}\text{O}$ [16] 2.39/0.43 1.40/0.34 1.48/0.41 1.89/0.63 1.55/0.53 1.53/0.50 1.34/0.39 1.98/0.53 1.81/0.57
MC-RANSAC [31] 4.79/1.27 3.76/1.00 4.39/1.40 - 6.60/1.74 4.67/1.19 4.76/1.08 6.49/1.73 7.72/2.19
MC-RANSAC + DPLR [31] 4.49/1.12 3.64/1.00 4.37/1.39 - 6.32/1.68 4.42/1.12 5.92/1.43 8.38/2.08 6.61/1.88
ORORA(Cen2018[17] 3.18/0.66 2.38/0.57 2.77/0.79 - 3.12/0.78 2.53/0.57 3.51/0.76 3.31/0.76 2.79/0.64
ORORA(Cen2019[17] 3.12/0.67 2.60/0.51 2.37/0.57 - 3.28/0.82 3.04/0.70 3.53/0.84 2.67/0.64 2.11/0.49
RINO(Cen2018) 1.64/0.46 1.25/0.30 0.91/0.39 1.15/0.43 1.42/0.46 2.08/1.08 2.28/0.53 2.35/0.80 2.28/0.47
RINO(Cen2019) 1.71/0.49 1.23/0.33 1.08/0.35 1.08/0.43 1.32/0.45 1.87/0.96 2.40/0.50 2.11/0.64 2.20/0.48
@@ IV-B Benchmark Results Within MulRan Dataset
We selected the best odometry or SLAM algorithms using different sensor modalities in large-scale environments as benchmarks. These methods include LOAM [7], SuMa [41], and the state-of-the-art scanning radar odometry algorithms introduced in Sec.II: CFEAR [14], SDPO [15], $\text{R}^{3}\text{O}$ [16], MC-RANSAC [31], and ORORA [17]. Specifically, CFEAR is evaluated in four configurations (CFEAR-1/2/3/3-s50) differing in solving strategy, keyframe settings, and robustness tuning. MC-RANSAC estimates motion with distortion compensation, while MC-RANSAC+DPLR further handles Doppler distortion. ORORA is tested with different feature extraction methods (Cen2018 and Cen2019) to assess the impact of feature quality.
@@
We conducted tests on nine sequences from the MulRan dataset: DCC01, DCC02, DCC03, KAIST01, KAIST02, KAIST03, Riverside01, Riverside02, and Riverside03. The Sejong sequences were excluded from evaluation, as they are designed for global localization rather than odometry, due to their large structural variations. The detailed results are shown in Tab.II. In the table, the red data points represent the minimum errors, while the blue data points correspond to the second smallest errors. It is recommended to view the table in color for improved clarity. It is important to note that, due to the unavailability of direct reproductions of these methods, the data presented in the table is taken from the reported results in publicly published papers. The table categorizes the aforementioned methods into LiDAR odometry, state-of-the-art radar odometry methods, and the baselines. The data is presented in the format (%/°/100m), where the first value represents the relative translation error, and the second value represents the relative rotation error.
@@
The proposed method, RINO, demonstrates relatively small relative translation and rotation errors across the majority of sequences, with particularly impressive performance in the DCC and KAIST sequences. RINO significantly improves upon the baseline, ORORA. Specifically, compared to ORORA with Cen2018 and Cen2019 keypoint extraction, RINO achieves average reductions of 1.43% and 1.73% in translation error, and 0.20°/100m and 0.13°/100m in rotation error in the DCC and KAIST sequences, respectively. In the DCC03 sequence, the error reduction reaches up to 1.46% in translation and 0.28°/100m in rotation. Moreover, when compared to LiDAR-based methods such as LOAM and SuMa, which offer higher resolution and suffer less from per-frame motion distortion, RINO has demonstrated superior performance. We conclude that RINO has achieved a level of radar odometry performance comparable to the state-of-the-art.
@@
When RINO is tested on the Riverside sequences, there is a notable decline in odometry accuracy compared to the DCC and KAIST sequences. The Riverside sequences, which were collected along a riverbank, represent a degenerate scenario that challenges odometry systems. Furthermore, the performance degradation of RINO is more pronounced in these sequences compared to other algorithms. A possible explanation for this lies in the Markovian assumption inherent in our pipeline. Both the Extended Kalman Filter (EKF) and the odometry based on global point cloud registration operate in an incremental fashion, where the current pose transformation estimate is solely dependent on the previous time step. This approach, to some extent, overlooks the broader structural information of point clouds in the local map, potentially leading to confusion in degenerate environments.
@@ IV-C1 Effectiveness of Adaptive Voting
We implemented a scanning radar odometry method, RO(GNC), using progressive non-convex optimization to solve pointcloud registration problem, which represents ORORA receiving sensor data via ROS (Robot Operating System). For comparison, we also implemented RO(AV) using adaptive voting to solve pointcloud registration, which represents the scanning radar branch in RINO. Both RO(GNC) and RO(AV) employ the improved keypoint extraction technique described in Sec. III-C, which differs from the code released by ORORA. Before feature matching, no voxel grid filtering is applied to the pointcloud in either method. This decision was based on our observations during the replication process, where we found that performing this step disrupted the timestamps assigned sequentially to each keypoint, thereby invalidating the motion distortion compensation. RO(GNC) performance is degraded compared to the test results presented in the ORORA paper (Tab.II). This degradation is hypothesized to arise from differences in parameter settings and the loss of precise azimuth encoding information in the raw scanning radar data. The comparison results between the two methods are shown in Tab.III. The odometry errors of RO(GNC) and RO(AV) are similar across multiple sequences. However, in subsequent runtime tests (Fig.10), we found that the running time of RO(AV) is generally shorter than that of RO(GNC). Therefore, we conclude that the adaptive voting method improves the system’s efficiency.
@@ IV-C2 Effectiveness of Adaptive Coupling
TABLE VI: Component-wise runtime for RIO(GNC) and RIO(AV). All times are reported in seconds [s].
Module RIO(GNC) RIO(AV)
[…]
Total 0.1458 0.1381
@@ IV-C4 Visualization
IV-D Runtime
We provide a detailed breakdown of the runtime for each module in RINO, including keypoint extraction and matching, IMU prediction, motion undistortion, pose estimation and fusion, and point cloud map publishing. Similar to ORORA, the majority of the system’s runtime is spent on keypoint extraction and matching. Notably, the runtime of the point cloud registration module is influenced by the number of keypoint correspondences, and its efficiency tends to degrade when the vehicle is moving slowly or remains stationary.
While a direct runtime comparison with ORORA is desirable, it is not presented due to differences in execution environments and data pipelines. Specifically, ORORA operates offline by reading pre-recorded data, whereas RINO simulates real-time operation via ROS message publishing, introducing additional system overhead. To provide a more meaningful performance comparison, we instead analyze the runtime of RIO(GNC)—which adopts the same robust optimization strategy as ORORA—and RIO(AV), which incorporates our proposed Adaptive Voting approach. This comparison offers insight into the computational efficiency improvements achieved by our enhancements.
As shown in Tab.VI, the average runtime of RINO is approximately 0.1381s per frame, ensuring real-time performance at the radar acquisition frequency of 4Hz. Furthermore, Fig.10 compares RIO(GNC) and RIO(AV), and shows that the adoption of the Adaptive Voting strategy results in an 11.5% improvement in the efficiency of pointcloud registration. In addition, the reduced variance observed in RIO(AV) indicates that the Adaptive Voting solver exhibits greater stability compared to the GNC-based approach.
IV-E Discussion about Estimation Uncertainty